OpenStudy (anonymous):

Select the ordered pair from the choices below that is a solution to the following system of equations: -7x + 6y = 11 8x − 3y = -1

OpenStudy (anonymous):

multiply 2nd equn. by 2 16x - 6y = -2 add this to 2nd equn. 9x = 9 x = 1 plug this into 2nd equn. 8 - 3y = -1 3y = 9 y=3 ordered pair is (1.3)

OpenStudy (anonymous):

(1)-7x+6y=11 (2)8x-3y=-1=> (x2)=> 16x-6y=-2 (1)+[(2)x2]=> -7x+16x+6y-6y=11-2=>9x=9=>x=1 (1) for x=1=>y=3 so (1.3)